# The Structure of Circle Graph States

### 09 Mar 2026 v1

@@ I.1 Results
I.2 Discussion
I.2.1 Graph states representation of general surface codes
@@
I.2.2 Closure of circle graph states under SLOCC
@@
I.2.3 Unifying the proofs of efficient simulation
The now two known proofs that MBQC on circle graph states is efficiently classically simulable stem from two distinct interpretation of circle graphs. First, bipartite circle graphs are the fundamental graphs of planar graphs, which is why they are connected to planar code states (see Section IV). Second, circle graphs can be interpreted as Eulerian tours of 4-regular multigraphs, leading to the connection with fermionic Gaussian states [36]. Finding unifying principles that explain when quantum systems are efficiently classically simulable is of interest (see, for example, [4]), and circle graph states may constitute a key toy example.
I.2.4 Geometric measure of entanglement of circle graph states
@@
I.2.5 Circle graphs in graph theory
@@
It is known that every graph of sufficiently large rank-width contains a vertex-minor isomorphic to any given circle graph [28]. Similarly, graphs of large rank-width are conjectured to contain pivot-minors isomorphic to any bipartite circle graph (4; see [54]).
Oum conjectures that graphs are well-quasi-ordered by the pivot-minor relation (1; see Question 6 in [55]) and by the vertex-minor relation (2; see [44]).
Geelen conjectures that MBQC is efficiently classically simulable on any vertex-minor-closed class of graphs (3; see also [48], Section 1.4]).
I.2.6 Quantum communication and quantum cryptography
@@
II Background and Definitions
II.1 Graph theory
@@
A graph $G$ is called bipartite if its vertex set can be split into two disjoint sets $A$ and $B$ such that every edge of $G$ connects a vertex in $A$ to a vertex in $B$. In simpler terms, the vertices of a bipartite graph may be colored into two colors such that no edge connects vertices of the same color. Simple graphs are bipartite if and only if they do not contain a cycle of odd length. (See the violet box of Figure 5 for an example of a graph that is not bipartite, and the orange box of the same figure for one that is.)
II.2 Graph states
@@
II.3 Local complementation and r-local complementation
@@
Local complementation with respect to a vertex $u$ of $G$ complements the subgraph induced by the neighborhood of $u$: Formally, this creates the graph $G\star u=G\Delta K_{N_{G}(u)}$, where $\Delta$ denotes the symmetric difference and $K_{N_{G}(u)}$ is the complete graph on the neighborhood of $u$. (Given two sets $A,B$, the symmetric difference of $A$ and $B$ is $A\Delta B\mathrel{\mathop{:}}=(A\cup B)\setminus(A\cap B)$.) See Figure 1 for an example.
@@
II.4 Vertex-minors and pivot-minors
@@ II.5 Circle graphs
Circle graphs are defined as follows, and examples are illustrated in Figure 1.
@@
A graph is a circle graph if and only if none of its vertex-minors is isomorphic to a graph in Figure 2, left.
@@
Bipartite circle graphs have chord diagrams whose chords are divided into two sets of chords that only intersect with chords from the other set and never with chords from the same set (see, e.g. Figure 4 on the right or Figure 5 on the bottom). Bipartite circle graphs are closed by pivoting and vertex deletion, in other words they are closed by the pivot-minor relation. Also, similar to circle graphs that are characterized by their excluded vertex-minors, bipartite circle graphs can be characterized by their excluded pivot-minors.
@@
A graph is a bipartite circle graph if and only if none of its pivot-minors is isomorphic to a graph in Figure 2, right.
@@
Figure 3: For an exemplary 99-vertex circle graph $C$ with a choice of independent set $K=\{1,2,3,4\}$, we show how to construct the corresponding tree $T$ and multigraph $H$. Violet arrow: From the chord representation of $C$, we remove all chords not corresponding to $K=\{1,2,3,4\}$. This yields a simpler diagram of four nonintersecting chords that divide the circle into five regions. In each of the five circle regions, we then draw a vertex and connect those pairs of vertices that correspond to pairs of neighboring regions by a tree edge. We label each of the four tree edges with the name of the chord that separates the neighboring circle regions. Green arrow: Adding all vertices $e\in V(G)\setminus K$ (that lie outside the independent set $K$) as additional edges to this tree $T$, creates fundamental cycles $F_{e}$ in a new multigraph $H$. As an example, the fundamental cycles $\color[rgb]{0.9296875,0.43359375,0.109375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9296875,0.43359375,0.109375}F_{5}$} and $\color[rgb]{0.76171875,0.9296875,0.109375}\definecolor[named]{pgfstrokecolor}{rgb}{0.76171875,0.9296875,0.109375}F_{6}$} are highlighted. Note that in the chord diagram of $C$, $e$ intersects a different chord $v$ if and only if $v$ is an edge of the fundamental cycle $F_{e}$ of $H$. In our example, $F_{5}$, containing edge $2$, corresponds to the 55 and the 22 chords intersecting the chord diagram of $C$ (and $F_{6}$, containing edges $1,2$, corresponds to 66 intersecting both $1,2$ in the chord diagram of $C$).
@@
The above construction for independent sets within circle graphs will be instrumental in proving the technical Lemma 2 of Section III.
II.6 Planar code states
@@
The qubits of a planar code state correspond to the edges of a planar multigraph. We allow two or more edges between two vertices (hence the denomination ”multigraph”). We also allow self-loops (i.e. edges with both ends being the same vertex), but we may ignore them as they correspond to isolated qubits. Given a planar multigraph $P$ with $n$ edges, the corresponding $n$-qubit planar code state is defined by the following generators of its $2^{n}$ element stabilizer group:
@@
Note that according to Euler’s theorem for planar graphs (|V||E|+|F|=2|V|-|E|+|F|=2, where $|V|$, $|E|$ and $|F|$ are the numbers of vertices, edges, and faces, respectively), the number of stabilizer generators described as above is equal to two plus the number of qubits, i.e., $n+2$. Therefore, two of these generators are redundant. For an illustration of a planar code state as it is typically presented, refer to Figure 4.
III Circle graphs are closed under r-local complementation
In this section we prove that the only graph states LU-equivalent to circle graph states are circle graph states themselves. To this end, we prove that circle graph states are closed by $r$-local complementation (see Section II.3).
@@ III Circle graphs are closed under r-local complementation
For an example of the construction that now follows, see Figure 3.
@@
We now apply both Lemma 1 and Lemma 2.
@@
Let $C$ be a circle graph, and $S$ be an independent $r$-incident multiset of $C$. According to Lemma 1, there exists an independent $r$-incident multiset $S^{\prime}$ such that
@@
Suppose by contradiction that $\textup{\text{supp}}(S^{\prime})$ is non-empty. Thanks to our application of Lemma 1, we know that then $\textup{\text{supp}}(S^{\prime})$ contains at least one vertex of degree two or higher, and that $\textup{\text{supp}}(S^{\prime})$ does not contain twins.
Thus we can apply Lemma 2: There exist two distinct vertices $a,b\in V(C)\setminus\textup{\text{supp}}(S^{\prime})$ such that $|N_{C}(a)\cap N_{C}(b)\cap\textup{\text{supp}}(S^{\prime})|=1$, implying that $S^{\prime}$ is not $r$-incident. Indeed, if $w\in\textup{\text{supp}}(S^{\prime})$ is the unique common neighbor of $a$ and $b$ in $\textup{\text{supp}}(S^{\prime})$, i.e., $N_{C}(a)\cap N_{C}(b)\cap\textup{\text{supp}}(S^{\prime})=\{w\}$ then $\sum_{u\in\bigcap_{v\in\{a,b\}}N_{C}(v)}S^{\prime}(u)=S^{\prime}(w)$, which is not a multiple of $2^{r-1}$, as $S^{\prime}(w)\in[1,2^{r-1}-1]$. (c.f. $k=0,\delta(k)=1$ in the definition of $r$-incidence.)
@@
IV Planar code states are bipartite circle graph states
The planar code states introduced in Section II.6 correspond to bipartite circle graph states introduced in Section II.5.
@@ IV Planar code states are bipartite circle graph states
This was proven in [62], but with some non-trivial restrictions, that translate to the corresponding bipartite circle graph not having any vertex of degree 1 or twins. For example, their proof does not work for the planar code state defined of a 2D grid (see Figure 4 for example).
@@
MBQC being efficiently classically simulable means here that there exists a classical polynomial time algorithm that approximately samples from the same output distribution as any sequence of (possibly adaptative) single-qubit measurements. Theorem 3 was first proved in [11] under the assumption that the measured qubits are connected, but this restriction was removed in [10]. This result then translates to bipartite circle graphs, as MBQC being efficiently classically simulable for a class of quantum states is invariant by local operations.
@@
To lift Corollary 2 to (possibly non-bipartite) circle graphs, we first need to reduce circle graphs to bipartite circle graphs.
@@
Figure 5: Every circle graph $C$ is a vertex-minor of some bipartite circle graph $B$. We show how to construct such a bipartite $B$ for the example where $C$ is a 5-cycle. 1. From a chord diagram of $C$, we read off (clockwise or counterclockwise) a double occurrence word (aebacbdcedaebacbdced) and draw a 4-regular multigraph with vertices $a,b,c,d,e$ and edges for all adjacent letters in this word (first and last letter are considered adjacent). 2. This 4-regular multigraph generally has a nonzero crossing number, but we can planarize it. Here, the crossing number is one and thus adding a single vertex $f$ creates a planar 4-regular multigraph. We then 2-color the faces of this graph in green and white. 3. Choose one face color (w.l.o.g. green). Draw a new graph with vertices for each green face and edges for each point where the green faces touch. 4. Select a spanning tree and draw a contour around it. 5. Rotate the edges of the spanning tree so that they connect to the contour. 6. Transform the contour into a circle, pulling the edges that lie outside the spanning tree into the circle. The result is a chord diagram of a bipartite circle graph.
@@
V Counting LU-equivalent graph states is $\mathsf{P}-$hard
@@
According to Theorem 1, for circle graph states, LC and LU-equivalence coincide, so the number of graph states LU-equivalent to a given circle graph state, is the same as the number of graph states LC-equivalent to it. Thus, counting the number of graph states LU-equivalent to a given graph state is $\mathsf{P}-$hard.
@@
VI Acknowledgments
The authors would like to thank Jim Geelen for many helpful discussions on circle graphs. Further, the authors would like to express their gratitude to Anna Pappa, Isaac D. Smith, Robert Raußendorf and Hans J. Briegel for useful discussions. This research was funded in part by the Austrian Science Fund (FWF) [SFB BeyondC F7102, DOI: 10.55776/F71; WIT9503323, DOI: 10.55776/WIT9503323]. For open access purposes, the authors have applied a CC BY public copyright license to any author-accepted manuscript version arising from this submission. This work was also supported by the European Union via the ERC Advanced Grant, QuantAI, (No. 101055129) and via the Quantum Internet Alliance (No. 101102140). The views and opinions expressed in this article are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council - neither the European Union nor the granting authority can be held responsible for them.
@@ Appendix A Rank-width of circle graphs
In order to give a lower bound on the rank-width of circle graphs, we first define rank-width and then lower bound it for a specific subset of circle graphs called comparability grids.
@@
Thus a comparability grid is an example of a comparability graph, i.e., a graph whose edges represent the comparable elements in some partially ordered set. Moreover, comparability grids are circle graphs; see Figure 6 for the $3\times 3$ comparability grid and one of its chord diagrams.
@@
Choose a row or column with as many vertices in $X$ as possible. By symmetry between the rows and columns (this symmetry may be most clear by referring to either Figure 6), we may assume that it is a column, say column $i$. Let $k$ be the number of vertices in column $i$ which are in $X$. Going for a contradiction, suppose that $X$ contains at least a third of the vertices of the graph. So at least a third of column $i$ is in $X$, and thus $k\geq n/3$.
@@
We restrict ourselves to considering only the $k$ rows whose $i$-th vertex is in $X$. We say that such a row $j$ is left-holed (respectively, right-holed) if it contains a vertex $v_{j}$ to the left (respectively, right) of column $i$ which is not in $X$. We call this vertex $v_{j}$ a left-hole (respectively, a right-hole). We claim that at most n/4n/{4} rows are left-holed. To see this, consider the square submatrix of the adjacency matrix whose rows are the left-holes $v_{j}$ and whose columns are the corresponding vertices $(i,j)\in X$ (see Figure 7). Since $v_{j}$ is adjacent to the $j$-th through $n$-th vertices in column $i$, this matrix is triangular with $1$s on the diagonal. Thus it has full rank. This proves the desired claim. A similar proof shows that at most n/4n/{4} rows are right-holed.
@@
Figure 8: Finding a balanced partition. The degree-3 vertex $v$ adjacent to the edge $e$ on the $Y_{e}$ side has three incident branches, each with four, three, and six leaves, respectively. In the left box, the branch facing towards $X_{e}$ has strictly less than one third of the leaves ($\nicefrac{{4}}{{13}}<\nicefrac{{1}}{{3}}$). So at least one of the two branches facing towards $Y_{e}$ has strictly more than one third of the leaves; here, the one with six. Labeling the first edge of this branch as $e$’ gives a more balanced partition shown in the right box.
@@
Now, if $Y_{e}$ contained more than two-thirds of the vertices, then we could find an edge $e^{\prime}$ corresponding to a more balanced partition $(X_{e^{\prime}},Y_{e^{\prime}})$ by moving closer to $Y_{e}$. For an example of this, see Figure 8.
Note that this is always possible because $T$ is a ternary tree. The degree-3 vertex $v$ adjacent to the edge $e$ on the $Y_{e}$ side has three incident branches partitioning the leaves of $T$. The branch facing towards $X_{e}$ has strictly less than one third of the leaves. So at least one of the two branches facing towards $Y_{e}$ has strictly more than one third of the leaves. Labeling the first edge of this branch as $e$’ gives a more balanced partition. So in fact $Y_{e}$ contains at most two thirds of the vertices and thus $X_{e}$ contains at least one third of the vertices (but also at most half since $Y_{e}$ is the larger side), i.e. $n^{2}/3\leq|X_{e}|\leq n^{2}/2$.
But since $r(X_{e})\leq n/4$, we can also apply Lemma 3 and conclude $|X_{e}|<n^{2}/3$, yielding a contradiction. ∎

### 28 Apr 2026 v2

@@ 1.1 Results
1.2 Discussion
1.2.1 Graph states representation of general surface codes
@@
1.2.2 Closure of circle graph states under SLOCC
@@
1.2.3 Unifying the proofs of efficient simulation
The now two known proofs that MBQC on circle graph states is efficiently classically simulable stem from two distinct interpretation of circle graphs. First, bipartite circle graphs are the fundamental graphs of planar graphs, which is why they are connected to planar code states (see Section˜4). Second, circle graphs can be interpreted as Eulerian tours of 4-regular multigraphs, leading to the connection with fermionic Gaussian states [36]. Finding unifying principles that explain when quantum systems are efficiently classically simulable is of interest (see, for example, [4]), and circle graph states may constitute a key toy example.
1.2.4 Geometric measure of entanglement of circle graph states
@@
1.2.5 Circle graphs in graph theory
@@
It is known that every graph of sufficiently large rank-width contains a vertex-minor isomorphic to any given circle graph [28]. Similarly, graphs of large rank-width are conjectured to contain pivot-minors isomorphic to any bipartite circle graph (˜4; see [54]).
Oum conjectures that graphs are well-quasi-ordered by the pivot-minor relation (˜1; see Question 6 in [55]) and by the vertex-minor relation (˜2; see [44]).
Geelen conjectures that MBQC is efficiently classically simulable on any vertex-minor-closed class of graphs (˜3; see also [48], Section 1.4]).
1.2.6 Quantum communication and quantum cryptography
@@
2 Background and Definitions
2.1 Graph theory
@@
A graph $G$ is called bipartite if its vertex set can be split into two disjoint sets $A$ and $B$ such that every edge of $G$ connects a vertex in $A$ to a vertex in $B$. In simpler terms, the vertices of a bipartite graph may be colored into two colors such that no edge connects vertices of the same color. Simple graphs are bipartite if and only if they do not contain a cycle of odd length. (See the violet box of Figure˜5 for an example of a graph that is not bipartite, and the orange box of the same figure for one that is.)
2.2 Graph states
@@
2.3 Local complementation and r-local complementation
@@
Local complementation with respect to a vertex $u$ of $G$ complements the subgraph induced by the neighborhood of $u$: Formally, this creates the graph $G\star u=G\Delta K_{N_{G}(u)}$, where $\Delta$ denotes the symmetric difference and $K_{N_{G}(u)}$ is the complete graph on the neighborhood of $u$. (Given two sets $A,B$, the symmetric difference of $A$ and $B$ is $A\Delta B\mathrel{\mathop{:}}=(A\cup B)\setminus(A\cap B)$.) See Figure˜1 for an example.
@@
2.4 Vertex-minors and pivot-minors
@@ 2.5 Circle graphs
Circle graphs are defined as follows, and examples are illustrated in Figure˜1.
@@
A graph is a circle graph if and only if none of its vertex-minors is isomorphic to a graph in Figure˜2, left.
@@
Bipartite circle graphs have chord diagrams whose chords are divided into two sets of chords that only intersect with chords from the other set and never with chords from the same set (see, e.g. Figure˜4 on the right or Figure˜5 on the bottom). Bipartite circle graphs are closed by pivoting and vertex deletion, in other words they are closed by the pivot-minor relation. Also, similar to circle graphs that are characterized by their excluded vertex-minors, bipartite circle graphs can be characterized by their excluded pivot-minors.
@@
A graph is a bipartite circle graph if and only if none of its pivot-minors is isomorphic to a graph in Figure˜2, right.
@@
Figure 3: For an exemplary 99-vertex circle graph $C$ with a choice of independent set $K=\{1,2,3,4\}$, we show how to construct the corresponding tree $T$ and multigraph $H$. Violet arrow: From the chord representation of $C$, we remove all chords not corresponding to $K=\{1,2,3,4\}$. This yields a simpler diagram of four nonintersecting chords that divide the circle into five regions. In each of the five circle regions, we then draw a vertex and connect those pairs of vertices that correspond to pairs of neighboring regions by a tree edge. We label each of the four tree edges with the name of the chord that separates the neighboring circle regions. Green arrow: Adding all vertices $e\in V(G)\setminus K$ (that lie outside the independent set $K$) as additional edges to this tree $T$, creates fundamental cycles $F_{e}$ in a new multigraph $H$. As an example, the fundamental cycles $\color[rgb]{0.9296875,0.43359375,0.109375}F_{5}$} and $\color[rgb]{0.76171875,0.9296875,0.109375}F_{6}$} are highlighted. Note that in the chord diagram of $C$, $e$ intersects a different chord $v$ if and only if $v$ is an edge of the fundamental cycle $F_{e}$ of $H$. In our example, $F_{5}$, containing edge $2$, corresponds to the 55 and the 22 chords intersecting the chord diagram of $C$ (and $F_{6}$, containing edges $1,2$, corresponds to 66 intersecting both $1,2$ in the chord diagram of $C$).
@@
The above construction for independent sets within circle graphs will be instrumental in proving the technical Lemma˜2 of Section˜3.
2.6 Planar code states
@@
The qubits of a planar code state correspond to the edges of a planar multigraph. We allow two or more edges between two vertices (hence the denomination “multigraph”). We also allow self-loops (i.e. edges with both ends being the same vertex), but we may ignore them as they correspond to isolated qubits. Given a planar multigraph $P$ with $n$ edges, the corresponding $n$-qubit planar code state is defined by the following generators of its $2^{n}$ element stabilizer group:
@@
Note that according to Euler’s theorem for planar graphs (|V||E|+|F|=2|V|-|E|+|F|=2, where $|V|$, $|E|$ and $|F|$ are the numbers of vertices, edges, and faces, respectively), the number of stabilizer generators described as above is equal to two plus the number of qubits, i.e., $n+2$. Therefore, two of these generators are redundant. For an illustration of a planar code state as it is typically presented, refer to Figure˜4.
3 Circle graphs are closed under r-local complementation
In this section we prove that the only graph states LU-equivalent to circle graph states are circle graph states themselves. To this end, we prove that circle graph states are closed by $r$-local complementation (see Section˜2.3).
@@ 3 Circle graphs are closed under r-local complementation
For an example of the construction that now follows, see Figure˜3.
@@
We now apply both Lemma˜1 and Lemma˜2.
@@
Let $C$ be a circle graph, and $S$ be an independent $r$-incident multiset of $C$. According to Lemma˜1, there exists an independent $r$-incident multiset $S^{\prime}$ such that
@@
Suppose by contradiction that $\textup{\text{supp}}(S^{\prime})$ is non-empty. Thanks to our application of Lemma˜1, we know that then $\textup{\text{supp}}(S^{\prime})$ contains at least one vertex of degree two or higher, and that $\textup{\text{supp}}(S^{\prime})$ does not contain twins.
Thus we can apply Lemma˜2: There exist two distinct vertices $a,b\in V(C)\setminus\textup{\text{supp}}(S^{\prime})$ such that $|N_{C}(a)\cap N_{C}(b)\cap\textup{\text{supp}}(S^{\prime})|=1$, implying that $S^{\prime}$ is not $r$-incident. Indeed, if $w\in\textup{\text{supp}}(S^{\prime})$ is the unique common neighbor of $a$ and $b$ in $\textup{\text{supp}}(S^{\prime})$, i.e., $N_{C}(a)\cap N_{C}(b)\cap\textup{\text{supp}}(S^{\prime})=\{w\}$ then $\sum_{u\in\bigcap_{v\in\{a,b\}}N_{C}(v)}S^{\prime}(u)=S^{\prime}(w)$, which is not a multiple of $2^{r-1}$, as $S^{\prime}(w)\in[1,2^{r-1}-1]$. (c.f. $k=0,\delta(k)=1$ in the definition of $r$-incidence.)
@@
4 Planar code states are bipartite circle graph states
The planar code states introduced in Section˜2.6 correspond to bipartite circle graph states introduced in Section˜2.5.
@@ 4 Planar code states are bipartite circle graph states
This was proven in [62], but with some non-trivial restrictions, that translate to the corresponding bipartite circle graph not having any vertex of degree 1 or twins. For example, their proof does not work for the planar code state defined of a 2D grid (see Figure˜4 for example).
@@
MBQC being efficiently classically simulable means here that there exists a classical polynomial time algorithm that approximately samples from the same output distribution as any sequence of (possibly adaptative) single-qubit measurements. Theorem˜3 was first proved in [11] under the assumption that the measured qubits are connected, but this restriction was removed in [10]. This result then translates to bipartite circle graphs, as MBQC being efficiently classically simulable for a class of quantum states is invariant by local operations.
@@
To lift Corollary˜2 to (possibly non-bipartite) circle graphs, we first need to reduce circle graphs to bipartite circle graphs.
@@
Figure 5: Every circle graph $C$ is a vertex-minor of some bipartite circle graph $B$. We construct such a $B$ for the example where $C$ is a 5-cycle: $1.$ From a chord diagram of $C$, we read off a double occurrence word (aebacbdcedaebacbdced) and draw a 4-regular multigraph with vertices $a,b,c,d,e$ and edges for all adjacent letters in this word, i.e., we contract the chords. $2.$ This 4-regular multigraph may have nonzero crossing number, but we can planarize it. Here, the crossing number is one and thus adding a single vertex $f$ yields a planar 4-regular multigraph. We then 2-color the faces of this graph in green and white. $3.$ Choose one face color (w.l.o.g. green). Draw a new graph with vertices for each green face and edges for each point where the green faces touch. $4.$ Select a spanning tree and draw a contour around it. $5.$ Rotate the edges of the spanning tree so that they connect to the contour. $6.$ Transform the contour into a circle, yielding a chord diagram of a bipartite circle graph $B$.
@@
5 Counting LU-equivalent graph states is #P-hard
@@
According to Theorem˜1, for circle graph states, LC and LU-equivalence coincide, so the number of graph states LU-equivalent to a given circle graph state, is the same as the number of graph states LC-equivalent to it. Thus, counting the number of graph states LU-equivalent to a given graph state is $\mathsf{P}-$hard.
@@
The authors would like to thank Jim Geelen for many helpful discussions on circle graphs. Further, the authors would like to express their gratitude to Anna Pappa, Isaac D. Smith, Robert Raußendorf and Hans J. Briegel for useful discussions. This research was funded in part by the Austrian Science Fund (FWF) [SFB BeyondC F7102, DOI: 10.55776/F71; WIT9503323, DOI: 10.55776/WIT9503323]. For open access purposes, the authors have applied a $C$ BY public copyright license to any author-accepted manuscript version arising from this submission. This work was also supported by the European Union via the ERC Advanced Grant, QuantAI, (No. 101055129) and via the Quantum Internet Alliance (No. 101102140). The views and opinions expressed in this article are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council – neither the European Union nor the granting authority can be held responsible for them.
@@ Appendix A Rank-width of circle graphs
Thus a comparability grid is an example of a comparability graph, i.e., a graph whose edges represent the comparable elements in some partially ordered set. Moreover, comparability grids are circle graphs; see Figure˜6 for the $3\times 3$ comparability grid and one of its chord diagrams.
@@
Choose a row or column with as many vertices in $X$ as possible. By symmetry between the rows and columns (this symmetry may be most clear by referring to either Figure˜6), we may assume that it is a column, say column $i$. Let $k$ be the number of vertices in column $i$ which are in $X$. Going for a contradiction, suppose that $X$ contains at least a third of the vertices of the graph. So at least a third of column $i$ is in $X$, and thus $k\geq n/3$.
@@
We restrict ourselves to considering only the $k$ rows whose $i$-th vertex is in $X$. We say that such a row $j$ is left-holed (respectively, right-holed) if it contains a vertex $v_{j}$ to the left (respectively, right) of column $i$ which is not in $X$. We call this vertex $v_{j}$ a left-hole (respectively, a right-hole). We claim that at most n/4n/{4} rows are left-holed. To see this, consider the square submatrix of the adjacency matrix whose rows are the left-holes $v_{j}$ and whose columns are the corresponding vertices $(i,j)\in X$ (see Figure˜7). Since $v_{j}$ is adjacent to the $j$-th through $n$-th vertices in column $i$, this matrix is triangular with $1$s on the diagonal. Thus it has full rank. This proves the desired claim. A similar proof shows that at most n/4n/{4} rows are right-holed.
@@
Figure 8: Finding a balanced partition. The degree-3 vertex $v$ adjacent to the edge $e$ on the $Y_{e}$ side has three incident branches, each with four, three, and six leaves, respectively. In the left box, the branch facing towards $X_{e}$ has strictly less than one third of the leaves ($\nicefrac{{4}}{{13}}<\nicefrac{{1}}{{3}}$). So at least one of the two branches facing towards $Y_{e}$ has strictly more than one third of the leaves; here, the one with six. Labeling the first edge of this branch as $e^{\prime}$ gives a more balanced partition shown in the right box.
@@
Now, if $Y_{e}$ contained more than two-thirds of the vertices, then we could find an edge $e^{\prime}$ corresponding to a more balanced partition $(X_{e^{\prime}},Y_{e^{\prime}})$ by moving closer to $Y_{e}$. For an example of this, see Figure˜8.
Note that this is always possible because $T$ is a ternary tree. The degree-3 vertex $v$ adjacent to the edge $e$ on the $Y_{e}$ side has three incident branches partitioning the leaves of $T$. The branch facing towards $X_{e}$ has strictly less than one third of the leaves. So at least one of the two branches facing towards $Y_{e}$ has strictly more than one third of the leaves. Labeling the first edge of this branch as $e^{\prime}$ gives a more balanced partition. So in fact $Y_{e}$ contains at most two thirds of the vertices and thus $X_{e}$ contains at least one third of the vertices (but also at most half since $Y_{e}$ is the larger side), i.e. $n^{2}/3\leq|X_{e}|\leq n^{2}/2$.
But since $r(X_{e})\leq n/4$, we can also apply Lemma˜3 and conclude $|X_{e}|<n^{2}/3$, yielding a contradiction. ∎